Angle P and angle Q are supplementary. Angle P measures 5x and angle Q measures 3x – 12. Find the measure of angle P.
Answer by checkley77(12569)

5X+3X-12=180
8X=180+12
8X=192
X=192/8
X=24
5*24=120 ANS. FOR ANGLE P
Q=3*24-12
Q=72-12
Q=60
PROOF:
120+60=180
180=180